bailey wants to make a scale drawing of his house.his house is 32 feet long. in the scale drawing,2inches will repersent 10 feet. which is the best estimate for the length of bailey's house in the scale drawing.
Answer by checkley75(3666)

10/2=32/x now cross multiply & we get
10x=2*32
10x=64
x=64/10
x=6.4 inches would be the scale drawing of his house.
:)